Ziggy bought x cans of red goo for y dollars a can, and z cans of blue goo for 3y dollars a can. If he bought twice as much blue goo as red goo, then in terms of y, what was the average cost, in dollars, per can of goo?
a) 3y/2 b) 7y/3 c) 3y – y d) 2y e) 6y

Answer by palanisamy(496)

Ziggy bought x cans of red goo for y dollars a can
Total cost of red goo = xy
Also,Ziggy bought z cans of blue goo for 3y dollars a can.
Total cost of blue goo = z*3y=3yz
He bought twice as much blue goo as red goo.
So, z=2x
Total number of cans = z+x= 2x+x = 3x
Total cost of all goo = xy+3yz=xy+3y*2x=xy+6xy=7xy
Average cost per can = (7xy)/(x+z)=7xy/(x+2x)=7xy/3x=$ 7y/3